RC meeting and Run Plan, Wednesday November 9th
From SBS wiki
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Current Status
- 4 Run Plan After Configuration Change
- 5 Notices in Effect for Shift Crew
- 6 Shifts and Long Term Issues
- 7 Planned Accesses to the Hall
- 8 Special Requests
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
- We were down following the sweep yesterday at 4:30pm due to accelerator issues. This delayed beginning beam checkout until end of swing.
- During owl we were delayed during functional ion chamber tests due to two unresponsive ion chambers (beam dump left and compton). After some discussion with MCC, SSG made two accesses to correct the issue and restore function to the ion chambers.
- Early during day shift, we worked with MCC to have work on the 3rd pass separator done. No beam during this time.
- We are in beam permit now and working through beam checkout procedures.
- Target polarization is ongoing.
- run time
- During day, the cooling jet alarms indicated loss of nitrogen flow to the beryllium windows. Our single replacement canister on-hand had a leak. We located a spare (thanks Hall D!) and Bill installed it, restoring the cooling jets. Jack Segal is aware of the issue and we expect a replacement dewar today.
- Status of the beam checkout and operations moving forward. Update from Simona and open discussion.
- Laser gui in alarm handler. Bill is working on it.
- Heater RTD signals are inconsistent. Currently running at lower temperature. Possible update from target team.
- Calibration constant at the current kinematic setting. Open discussion.
- Current Polarization: 10.2mV -> 12.2mV (relative, upstream, calibration constant = 1, last two NMRs, growing!)
- If holding field is changed for any reason, do NOT perform an NMR measurement.
- Please be mindful of necessary screenshots and frequency detailed in target procedure.
- Hourly screenshots are necessary for the laser gui.
- Update from target team.
Any additional subsystem updates?
- Complete beam checkout.
- Take EPR opportunistically
- Move to production data as soon as polarization at acceptable level. (See current run plan procedure)
- Take reference cell data opportunistically.
Experimental progress (for MCC planning meeting)
- 6.1% of accumulated charge for production data (13.55 Coulombs out of 232)
- 32% Experimental run time (45 out of 141 days)
- Long term "to do" items; not urgent, some can be done when 3He target not available
- Luminosity scans with GEMs. See procedure.
- Opportunistic BCM calibration (higher current) before pass change?
- INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run
- Work and testing on LV supplies was performed during opportunistic access. See log.
- Misc. GEM commissioning items: https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29
Run Plan After Configuration Change
Continue to update 'standard' run plan for 3-pass data taking as necessary.
- GEM luminosity studies could be performed between beam restoration and target polarization. See procedure.
- NMR measurement every 2-3 hours. See procedure.
- Restore beam. One time process led by Simona before production.
- Optics and H2 Ref Cell (Sometime during run)
- Production Run Plan
- 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=0, all others -1. Take four runs that should each last for one hour. (Check ps5 with Alex)
- 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization.
- 3.) Return to step 1.)
While on shift:
- Keep an eye out for drifting BPM positions
- Log all DAQ crashes in the excel worksheet along with the reason for the crash
- If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: procedure, If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
- Cross check the rate with what is expected: log
- Cross check with the expected ion chamber readings: log
Notices in Effect for Shift Crew
All SL shifts are on standby today
Be sure to follow these procedures:
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
Shifts and Long Term Issues
- Shift Openings: We're looking good for now! Schedule
Planned Accesses to the Hall
In restricted Access. Survey and alignment ongoing.
Nothing scheduled at present.